Problem 3 Determine the y parameters of the circuits below. 1Ω 30Ω 1Ω 30Ω 1Ω 30 Ω 2Ω 40 Ω 40 Ω ΑΛΜ 1Ω 3Ω 40 Ω 4Ω 40 Ω 30 Ω
Answer #1

I 12 mtē + Izo z 10 V Looky Intiz Cleopa Ve le Iz Apply KVL in loop 1 V₂ = It Iz Apply KVL in loop 2 - V₂ + Iz X 1 + (I, + I₂I = Vi - V₂ us 90 Iz =-V2 V2 + V₂ 90 45 Y = 45 - 26 - 2 us v IS AN 22 B Iz Arms + + o) 332 V2 Vi {lon 2 a ܝܐ 2 2 uz Apply KCLЧОЛ- 402 + I Vz u 1 1 2 I va your 402 It Izor sm I Iz Apply KUL - V₂ + 40 I, + 40 (I + I₂) + 401, = 0 Vi = 1201, +40 I, ApplyV2 320 Is = 3 Vi 320 Iz = -V 3 V₂ + 320 320 Y = 3 320 - 1 320 -L 320 3 320 v

All answer of given circuits is shown above.